# Mamba®: Vision Mamba ALSO Needs Registers

### v1

@@ Abstract
Similar to Vision Transformers, this paper identifies artifacts also present within the feature maps of Vision Mamba. These artifacts, corresponding to high-norm tokens emerging in low-information background areas of images, appear much more severe in Vision Mamba—they exist prevalently even with the tiny-sized model and activate extensively across background regions. To mitigate this issue, we follow the prior solution of introducing register tokens into Vision Mamba. To better cope with Mamba blocks’ uni-directional inference paradigm, two key modifications are introduced: 1) evenly inserting registers throughout the input token sequence, and 2) recycling registers for final decision predictions. We term this new architecture Mamba®. Qualitative observations suggest, compared to vanilla Vision Mamba, Mamba®’s feature maps appear cleaner and more focused on semantically meaningful regions. Quantitatively, Mamba® attains stronger performance and scales better. For example, on the ImageNet benchmark, our Mamba®-B attains 82.9% accuracy, significantly outperforming Vim-B’s 81.8%; furthermore, we provide the first successful scaling to the large model size (i.e., with 341M parameters), attaining a competitive accuracy of 83.2% (84.5% if finetuned with $384\times 384$ inputs). Additional validation on the downstream semantic segmentation task also supports Mamba®’s efficacy.
The project page is https://wangf3014.github.io/mambar-page/.
@@ 1 Introduction
Recent advances in State Space Models (SSMs) have showcased their considerable potential in sequence modeling tasks. In contrast to Transformers’ quadratic computational complexity with respect to sequence lengths, SSMs operate with linear computational complexity, offering significant efficiency advantages in managing extended sequences. One exemplary instantiation of SSMs is the Mamba architecture (Gu & Dao, 2023), which employs selective scan techniques alongside a suite of hardware-optimized designs. This innovation facilitates the efficient training and inference of recurrent models with linear computational complexity and memory overhead. Furthermore, a comprehensive body of recent research (Gu & Dao, 2023; Behrouz et al., 2024; Lieber et al., 2024) substantiates that the Mamba architecture is able to achieve competitive performance levels, on par with Transformers, particularly in processing natural language and audio.
Furthermore, the Mamba architecture has also been successfully extended to a variety of visual tasks (Zhu et al., 2024; Liu et al., 2024b; Li et al., 2024; Hu et al., 2024). The motivation for this expansion mainly arises from the computational challenges presented by processing high-resolution images and videos. These data types often lead to long input sequences that conventional models struggle to handle effectively or efficiently—e.g., for long-length input, traditional models such as Convolutional Neural Networks (CNNs) suffer from relatively small receptive fields, and Vision Transformers (ViTs) contend with high computational and memory costs. Yet, Vision Mamba (Vim) architectures have shown the potential to mitigate these limitations—recent works demonstrate that they not only manage computational and memory demands more efficiently but also deliver strong performance across a variety of generic visual tasks, including classification, segmentation, and image generation (Zhu et al., 2024; Liu et al., 2024a; Hu et al., 2024).
Despite the competitive benchmark performance, our observations reveal that Mamba’s internal modeling exhibits significant issues when processing visual inputs. This issue is similar to the prior observation of ViTs (Darcet et al., 2024), where some outlier tokens located in the less semantic background unexpectedly contain rich global information (showing as high attention scores in the feature map). These unusual feature activations are termed artifacts. In this work, we reveal that this artifact issue not only exists but actually is considerably more severe in Vision Mamba. For example, the artifacts are clearly visible in the feature maps of Vim (Zhu et al., 2024), as illustrated in the 2nd and 4th columns of Figure 1, where activations encompass not just semantically significant content, but also extend to expansive yet minimally informative background regions. Furthermore, as quantitatively confirmed in Section 3.2, these artifact tokens are widely present in different sizes of Vision Mamba models and possess rich global information.
@@
Figure 1: Feature maps of vanilla Vision Mamba (Vim) (Zhu et al., 2024) and our Mamba®. It shows that massive artifacts appear in Vim’s feature map, making the model difficult to attend to visually meaningful content within the image. In contrast, our model exhibits much cleaner feature activations, showcasing the significant efficacy of our enhanced architectural design.
In this paper, building upon a previous solution (Darcet et al., 2024), we introduce a straightforward yet effective architectural refinement to Vision Mamba by appending registers—new, input-independent tokens—to the token sequence. Unlike (Darcet et al., 2024) which only appends several register tokens at one end of the input layer, we 1) insert register tokens evenly throughout the token sequence; and 2) at the end of the Vision Mamba, concatenate the register tokens to form a comprehensive image representation for the final prediction. We name this enhanced architecture Mamba®.
@@
Empirically, Mamba® showcases advantages on two fronts. Qualitatively, as evidenced by the cleaner feature maps displayed in the 3rd and 6th columns of Figure 1, Mamba® significantly suppresses artifacts, with responses now more focused on visually meaningful content. Meanwhile, as visualized in Figure 6, registers can well capture object-related semantic information for building high-quality image representations. Quantitatively, the improvements in benchmarks are equally compelling. For example, Mamba®-Base achieves an accuracy of 82.9% on ImageNet, notably outperforming the 81.8% accuracy of Vim-Base, which is a vanilla Vision Mamba architecture. Furthermore, Mamba® successfully expands the scaling capabilities of Vision Mamba—whereas previous models were limited to configurations no larger than 90M parameters (Zhu et al., 2024; Liu et al., 2024b; Yang et al., 2024), Mamba® can be effectively trained with up to 341M parameters, reaching an impressive 83.2% accuracy on ImageNet; this accuracy can be further boosted to 84.5% by enlarging the image input size to $384\times 384$. For the ADE20k (Zhou et al., 2019) semantic segmentation benchmark, our Mamba® attains a 49.1% evaluation mIoU, which significantly outperforms Vim’s best result of 44.9% mIoU (Zhu et al., 2024).
@@ 2 Related Work
Generic Visual Backbone Architectures. Modern computer vision predominantly relies on two types of backbone architectures: CNNs that excel in extracting hierarchical features and ViTs that are effective in modeling long-range dependencies. Since the advent of CNNs (LeCun et al., 1998), their structure and scale have undergone a series of significant innovations in recent decades (Krizhevsky et al., 2012; Simonyan & Zisserman, 2015; He et al., 2016; Huang et al., 2017; Tan & Le, 2019; Liu et al., 2022). Unlike CNNs that build spatial dependencies through convolutional operations, ViTs (Dosovitskiy et al., 2021) attain a global receptive field by utilizing the self-attention mechanism (Vaswani et al., 2017), leading to state-of-the-art performance in a series of downstream visual tasks. Based on this architecture, extensive research has been dedicated to improving its model design (Yuan et al., 2021; Chen et al., 2021a; Liu et al., 2021), enhancing training strategy (Touvron et al., 2021, 2022), and advancing self-supervised pretraining frameworks (Chen et al., 2021b; Caron et al., 2021; Bao et al., 2022; He et al., 2022).
State Space Models. The concept of State Space Models (SSMs) can be dated back to the 1960s in control systems where it was used to process continuous inputs (Kalman, 1960). With advancements in discretization strategies (Tallec & Ollivier, 2018; Gu et al., 2020; Nguyen et al., 2022; Gu et al., 2023), SSMs have recently been introduced into the field of deep learning, modeling sequential information such as language and speech (Gu et al., 2022, 2021; Smith et al., 2022). Broadly defined, SSMs can refer to any recurrent models with a latent state such as Hidden Markov Models (Eddy, 1996), RNNs, and the variant architectures such as Linear Attention (Katharopoulos et al., 2020), RetNet (Sun et al., 2023), and RWKV (Peng et al., 2023). More recently, Gu & Dao (2023) introduced a selective SSM block, namely Mamba, that incorporates structured SSMs with hardware-aware state expansion, leading to a highly efficient recurrent architecture that is competitive to Transformer.
Mamba Models in Vision. Building upon the Mamba block, a series of follow-up studies have explored the application of SSMs in computer vision. For example, Zhu et al. (2024) propose a straightforward vision Mamba model by sequentially stacking the Mamba blocks, attaining superior performance than Vision Transformers (Dosovitskiy et al., 2021; Touvron et al., 2021) in both tiny and small model sizes. Liu et al. (2024b) presents a hybrid architecture that combines Mamba with 2D convolution, showcasing significant results in a series of vision tasks. The study of Mamba-based architectures is continuously explored (Lieber et al., 2024; Li et al., 2024; Liu et al., 2024a).
@@ 3.1 Mamba Preliminaries
where ${\bm{A}}\in\mathbb{C}^{N\times N}$, ${\bm{B}}\in\mathbb{C}^{N}$, ${\bm{C}}\in\mathbb{C}^{N}$, and $D\in\mathbb{C}^{1}$ denote the weighting parameters.
By discretizing this ordinary differential equation group, the continuous-time SSMs can be integrated to process discrete inputs such as language, speech, and image pixels. To this end, the model can be solved by an analytic solution
[…]
and then approximated by Zero-Order Hold (Gu & Dao, 2023), leading to a discrete model as follows:
@@
However, the Structural State Space Models’ nature of Linear Time-Invariance significantly limits its capacity to fit contextual information, making it difficult to scale up and achieve performance comparable to Transformers. The Selective State Space Model, also known as Mamba or S6 (Gu & Dao, 2023), improves it by introducing input-dependent parameters ${\bm{B}}={\bm{S}}_{B}(x)$, ${\bm{C}}={\bm{S}}_{C}(x)$, and $\Delta={\bm{S}}_{\Delta}(x)$, leading to a time-varying system that can model more complex inputs. Notably, with associative scan algorithms (Martin & Cundy, 2018; Smith et al., 2022), the Mamba module can be trained and inferred efficiently via parallel computing, with detailed mathematical derivations elaborated in (Gu & Dao, 2023).
@@
To adapt Mamba for visual tasks, images are first processed into sequential inputs through patch embedding as in ViT. However, the standard Mamba is a unidirectional model where each token in the sequence can only access information from preceding tokens. This characteristic, while working well with 1-D language signals, significantly constrains the model’s capacity to gather contextual information inherently from 2-D visual signals. To overcome this limitation, a common solution is to reconfigure Mamba blocks for bidirectional scanning. Specifically, the sequence is scanned once from start to end and again from end to start, with the outputs from both scans subsequently averaged to obtain a comprehensive representation (Zhu et al., 2024). We follow this scanning design in all subsequent experiments.
3.2 Feature Artifacts of Vision Mamba
In ViT, an interpretable feature map can be obtained by visualizing the activation scores in their self-attention blocks. Ideally, under appropriate pre-training paradigms, these feature maps are expected to display high attention scores in the informative foreground regions of images and relatively low scores in less semantic background areas. Nevertheless, a considerable amount of outliers often appear in these feature maps, which position-wise correspond to low-information background regions yet exhibit anomalously high attention scores. A recent study (Darcet et al., 2024) has termed these outliers as feature artifacts. Specifically, this study reveals that the artifact tokens always possess high normalization values and, during inference, they tend to discard their local information in favor of retaining global features, thereby ‘compromising’ the quality of the feature map.
This work identifies a similar issue in Vision Mamba models. First, by computing the $\ell_{2}$ distances between vanilla Vision Mamba’s global and local outputs, we observe a considerable amount of activations in background areas (shown in Figure 1). Further analysis of their normalization reveals that these background activations also exhibit high normalization values, akin to the artifacts observed in ViTs. For example, by visualizing the $\ell_{2}$ normalization of vanilla Vision Mamba’s local outputs in Figure 2, we can observe a significant presence of high-norm tokens in the background, even blurring the distinction between foreground and background regions. Quantitatively, we plot the norm distributions of vanilla Vision Mamba in Figure 3(a), where it clearly displays a number of outliers with high normalization, confirming consistency with previous findings in ViTs as discussed in (Darcet et al., 2024).
Furthermore, it is equally noteworthy that these artifacts in Vision Mamba function similarly to those in ViTs in retaining global representations (Darcet et al., 2024). As reported in Table 4, the vanilla Vision Mamba model can obtain 81.0% ImageNet accuracy by merely using the average of top 5% high norm tokens as a global feature, which is only 0.1% lower than that of pooling all local tokens. Increasing this threshold to top 10% or 20% high norm tokens further enables the model to match the accuracy of using global pooling. In contrast, relying on the remaining 80% of relatively low-norm tokens results in a performance drop to 79.3%.
(a) Norm distribution of vanilla Vision Mamba
(b) Norm distribution of Mamba®
@@ 3.2 Feature Artifacts of Vision Mamba
Figure 3: Distributions of $\ell_{2}$ normalization values of local outputs across different layers. It quantitatively shows that our Mamba®effectively reduces the number of high-norm outliers.
@@
Figure 4: Normalization distribution across different sizes of Vision Mamba.
@@ 3.3 Mamba®: Vision Mamba with Registers
Following the solution of removing artifacts in ViT  (Darcet et al., 2024), we propose to address this issue by introducing register tokens into Vision Mamba. We term our new architecture Mamba®. Unlike the previous method (Darcet et al., 2024) which only appends register tokens at one end of the input sequence, we hypothesize that by distributing the register tokens more densely throughout the sequence, our method can 1) better address the more pervasive artifact issue that is unique to vision mamba; and 2) helps capture global representation that is often missed in vision mamba due to its uni-directional nature. The framework of Mamba® is illustrated in Figure 5. Overall, we follow the backbone architecture of vanilla Vision Mamba (Vim) (Zhu et al., 2024), where the input image is first decomposed into a sequence of non-overlapping patches and then fed into a stack of bi-directional Mamba blocks. Based on this plain architecture, we make the following two simple yet very effective modifications to build our Mamba®.

### v2

@@ Abstract
Similar to Vision Transformers, this paper identifies artifacts also present within the feature maps of Vision Mamba. These artifacts, corresponding to high-norm tokens emerging in low-information background areas of images, appear much more severe in Vision Mamba—they exist prevalently even with the tiny-sized model and activate extensively across background regions. To mitigate this issue, we follow the prior solution of introducing register tokens into Vision Mamba. To better cope with Mamba blocks’ uni-directional inference paradigm, two key modifications are introduced: 1) evenly inserting registers throughout the input token sequence, and 2) recycling registers for final decision predictions. We term this new architecture Mamba®. Qualitative observations suggest, compared to vanilla Vision Mamba, Mamba®’s feature maps appear cleaner and more focused on semantically meaningful regions. Quantitatively, Mamba®attains stronger performance and scales better. For example, on the ImageNet benchmark, our Mamba®-B attains 83.0% accuracy, significantly outperforming Vim-B’s 81.8%; furthermore, we provide the first successful scaling to the large model size with 341M parameters, attaining competitive accuracies of 83.6% and 84.5% for 224$\times$224 and 384$\times$384 inputs, respectively. Additional validation on the downstream semantic segmentation task also supports Mamba®’s efficacy. Code is available at https://github.com/wangf3014/Mamba-Reg.
@@ 1 Introduction
Figure 1: Feature maps of vanilla Vision Mamba (Vim) [54] and our Mamba®. It shows that massive artifacts appear in Vim’s feature map, making the model difficult to attend to visually meaningful content within the image. In contrast, our model exhibits much cleaner feature activations, showcasing the significant efficacy of our enhanced architectural design.
Recent advances in State Space Models (SSMs) have showcased their considerable potential in sequence modeling tasks. In contrast to Transformers’ quadratic computational complexity with respect to sequence lengths, SSMs operate with linear computational complexity, offering significant efficiency advantages in managing extended sequences. One exemplary instantiation of SSMs is the Mamba architecture [9], which employs selective scan techniques alongside a suite of hardware-optimized designs. This innovation facilitates the efficient training and inference of recurrent models with linear computational complexity and memory overhead. Furthermore, a comprehensive body of recent research [9, 2, 25] substantiates that the Mamba architecture is able to achieve competitive performance levels, on par with Transformers, particularly in processing natural language and audio.
Furthermore, the Mamba architecture has also been successfully extended to a variety of visual tasks [54, 27, 24, 17]. The motivation for this expansion mainly arises from the computational challenges presented by processing high-resolution images and videos. These data types often lead to long input sequences that conventional models struggle to handle effectively or efficiently—e.g., for long-length input, traditional models such as Convolutional Neural Networks (CNNs) suffer from relatively small receptive fields, and Vision Transformers (ViTs) contend with high computational and memory costs. Yet, Vision Mamba (Vim) architectures have shown the potential to mitigate these limitations—recent works demonstrate that they not only manage computational and memory demands more efficiently but also deliver strong performance across a variety of generic visual tasks, including classification, segmentation, and image generation [54, 26, 17].
Despite the competitive benchmark performance, our observations reveal that Mamba’s internal modeling exhibits significant issues when processing visual inputs. This issue is similar to the prior observation of ViTs [6], where some outlier tokens located in the less semantic background unexpectedly contain rich global information (showing as high attention scores in the feature map). These unusual feature activations are typically termed artifacts in the existing studies [6]. In this work, we reveal that this artifact issue not only exists but actually is considerably more severe in Vision Mamba. For example, the artifacts are clearly visible in the feature maps of Vim [54], as illustrated in the 2nd and 4th columns of Figure 1, where activations encompass not just semantically significant content, but also extend to expansive yet minimally informative background regions. Furthermore, as quantitatively confirmed in Section 3.2, these artifact tokens are widely present in different sizes of Vision Mamba models and possess rich global information. Building upon a previous solution [6], we introduce a straightforward yet effective architectural refinement to Vision Mamba by appending registers—new, input-independent tokens—to the token sequence. Unlike [6] which only appends several register tokens at one end of the input layer, we 1) insert register tokens evenly throughout the token sequence; and 2) at the end of the Vision Mamba, concatenate the register tokens to form a comprehensive image representation for the final prediction. We name this enhanced architecture Mamba®.
@@
Empirically, Mamba® showcases advantages on two fronts. Qualitatively, as evidenced by the cleaner feature maps displayed in Figure 1, Mamba® significantly suppresses artifacts, with responses now more focused on visually meaningful content. Meanwhile, as visualized in Figure 6, registers can well capture object-related semantic information for building high-quality image representations. Quantitatively, the improvements in benchmarks are equally compelling. For example, Mamba®-Base achieves an accuracy of 83.0% on ImageNet, notably outperforming the 81.8% accuracy of the vanilla Vim-Base. Furthermore, Mamba® successfully expands the scaling capabilities of Vision Mamba—whereas previous models were limited to configurations no larger than 90M parameters [54, 27, 50], Mamba® can be effectively trained with up to 341M parameters, reaching an impressive 83.6% accuracy on ImageNet; this accuracy can be further boosted to 84.5% by enlarging the image input size to $384\times 384$. For the ADE20k [52] semantic segmentation benchmark, our Mamba® attains a 49.1% evaluation mIoU, which significantly outperforms Vim’s best result of 44.9% mIoU [54].
(a) Norm distribution of vanilla Vision Mamba
(b) Norm distribution of Mamba®
Figure 3: Distributions of $\ell_{2}$ normalization values of local outputs across different layers. Compared with the vanilla Vision Mamba model, it quantitatively shows that our Mamba® effectively reduces the number of high-norm outliers.
@@ 2 Related Work
Generic Visual Backbone Architectures. Modern computer vision predominantly relies on two types of backbone architectures: CNNs that excel in extracting hierarchical features and ViTs that are effective in modeling long-range dependencies. Since the advent of CNNs [23], their structure and scale have undergone a series of significant innovations in recent decades [22, 37, 15, 18, 41, 29]. Unlike CNNs that build spatial dependencies through convolutional operations, ViTs [8] attain a global receptive field by utilizing the self-attention mechanism [44], leading to state-of-the-art performance in a series of downstream visual tasks. Based on this architecture, extensive research has been dedicated to improving its model design [51, 4, 28], enhancing training strategy [42, 43], and advancing self-supervised pretraining and inference frameworks [5, 3, 1, 16, 34, 53, 45].
State Space Models. The concept of State Space Models (SSMs) can be dated back to the 1960s in control systems where it was used to process continuous inputs [20]. With advancements in discretization strategies [40, 10, 32, 13], SSMs have recently been introduced into the field of deep learning, modeling sequential information such as language and speech [12, 11, 38]. Broadly defined, SSMs can refer to any recurrent models with a latent state such as RNNs and the variant architectures such as Linear Attention [21], RetNet [39], and RWKV [33]. More recently, [9] introduced a selective SSM block, namely Mamba, that incorporates structured SSMs with hardware-aware state expansion, leading to a highly efficient recurrent architecture that is competitive to Transformer.
Mamba Models in Vision. Building upon the Mamba block, a series of follow-up studies have explored the application of SSMs in computer vision. For example, [54] propose a straightforward vision Mamba model by sequentially stacking the Mamba blocks, attaining superior performance than Vision Transformers [8, 42] in both tiny and small model sizes. [27] presents a hierarchical Mamba architecture, showcasing significant results in a series of vision tasks. The study of Mamba-based architectures is continuously explored [14, 46, 47, 35, 36, 25, 48, 24, 26].
@@ 3.1 Mamba Preliminaries
where ${\bm{A}}\in\mathbb{R}^{N\times N}$, ${\bm{B}}\in\mathbb{R}^{N}$, ${\bm{C}}\in\mathbb{R}^{N}$, and $D\in\mathbb{R}^{1}$ denote the weighting parameters.
By discretizing this ordinary differential equation group, the continuous-time SSMs can be integrated to process discrete inputs such as language, speech, and image pixels. To this end, the model can be solved by an analytic solution and then approximated by Zero-Order Hold [9], leading to a discrete model:
@@
However, the Structural State Space Models’ nature of Linear Time-Invariance significantly limits its capacity to fit contextual information, making it difficult to scale up and achieve performance comparable to Transformers. The Selective State Space Model, also known as Mamba or S6 [9], improves it by introducing input-dependent parameters ${\bm{B}}={\bm{S}}_{B}(x)$, ${\bm{C}}={\bm{S}}_{C}(x)$, and $\Delta={\bm{S}}_{\Delta}(x)$, leading to a time-varying system that can model more complex inputs. Notably, with associative scan algorithms [31, 38], the Mamba module can be trained and inferred efficiently via parallel computing, with detailed mathematical derivations elaborated in [9].
@@
To adapt Mamba for visual tasks, images are first processed into sequential inputs through patch embedding as in ViT. However, the standard Mamba is a unidirectional model where each token in the sequence can only access information from preceding tokens. This characteristic, while working well with 1-D language signals, significantly constrains the model’s capacity to gather contextual information inherently from 2-D visual signals. To overcome this limitation, a common solution is to reconfigure Mamba blocks for bidirectional scanning. Specifically, the sequence is scanned once from start to end and again from end to start, with the outputs from both scans subsequently averaged to obtain a comprehensive representation [54]. We follow this scanning design in all subsequent experiments.
@@
Figure 4: Distributions of $\ell_{2}$ normalization across different sizes of Vision Mamba models. It shows that high-norm outliers widely appear in different sizes of Mamba models.
@@
3.2 Feature Artifacts of Vision Mamba
In ViT, an interpretable feature map can be obtained by visualizing the activation scores in their self-attention blocks. Ideally, under appropriate pre-training paradigms, these feature maps are expected to display high attention scores in the informative foreground regions of images and relatively low scores in less semantic background areas. Nevertheless, a considerable amount of outliers often appear in these feature maps, which position-wise correspond to low-information background regions yet exhibit anomalously high attention scores. A recent study [6] has termed these outliers as feature artifacts. Specifically, this study reveals that the artifact tokens always possess high normalization values and, during inference, they tend to discard their local information in favor of retaining global features, thereby “compromising” the quality of the feature map.
This work identifies a similar issue in Vision Mamba models. First, by computing the $\ell_{2}$ distances between vanilla Vision Mamba’s global and local outputs, we observe a considerable amount of activations in background areas (shown in Figure 1). Further analysis of their normalization reveals that these background activations also exhibit high normalization values, akin to the artifacts observed in ViTs. For example, by visualizing the $\ell_{2}$ normalization of vanilla Vision Mamba’s local outputs in Figure 2, we can observe a significant presence of high-norm tokens in the background, even blurring the distinction between foreground and background regions. Quantitatively, we plot the norm distributions of vanilla Vision Mamba in Figure 3(a), where it clearly displays a number of outliers with high normalization, confirming consistency with previous findings in ViTs as discussed in [6].
Furthermore, it is equally noteworthy that these artifacts in Vision Mamba function similarly to those in ViTs in retaining global representations [6]. As reported in Table 1, the vanilla Vision Mamba model can obtain 81.0% ImageNet accuracy by merely using the average of top 5% high norm tokens as a global feature, which is only 0.1% lower than that of pooling all local tokens. Increasing this threshold to top 10% or 20% high norm tokens further enables the model to match the accuracy of using global pooling. In contrast, relying on the remaining 80% of relatively low-norm tokens results in a performance drop to 79.3%.
@@ 3.3 Mamba®: Vision Mamba with Registers
Following the solution of removing artifacts in ViT  [6], we propose to address this issue by introducing register tokens into Vision Mamba. We term our new architecture Mamba®. Unlike the previous method [6] which only appends register tokens at one end of the input sequence, we hypothesize that by distributing the register tokens more densely throughout the sequence, our method can 1) better address the more pervasive artifact issue that is unique to vision mamba; and 2) helps capture global representation that is often missed in vision mamba due to its uni-directional nature. The framework of Mamba® is illustrated in Figure 5. Overall, we follow the backbone architecture of vanilla Vision Mamba (Vim) [54], where the input image is first decomposed into a sequence of non-overlapping patches and then fed into a stack of bi-directional Mamba blocks. Based on this plain architecture, we make the following two simple yet very effective modifications to build our Mamba®.